OpenStudy (anonymous):

which of the following CANNOT be exprssed as the sum of three consecutive integers? a)18 b)24 c)33 d)36

Vocaloid (vocaloid):

hm... we can write the sum of three consecutive integers as n + (n + 1) + (n+2) = 3n + 3 = (the sum) we know that n must be an integer, so 3n + 3 must be divisible by 3

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

n-1,n,n+1 n-1+n+n+1=3n each term is divisible by 3

OpenStudy (anonymous):

correction middle term is ,divide by 3

OpenStudy (anonymous):

s the answer 18?

OpenStudy (anonymous):

e.g.,24/3=8 middle term is 8 8-1,8,8+1or 7,8,9 are the required terms

OpenStudy (anonymous):

sum is divisible by 3 for 18 18/3=6 6-1,6,6+1 or 5,6,7

OpenStudy (anonymous):

each term means all your choices are divisible by 3

OpenStudy (anonymous):

answer is 18?

OpenStudy (anonymous):

5+6+7=18 there is no term .

OpenStudy (loser66):

To me all of them can be expressed as sum of 3 consecutive numbers.

OpenStudy (anonymous):

you are absolutely correct.

OpenStudy (loser66):

3| 18 =6, hence middle one is 6, so the numbers is 5 , 6 ,7 3|24 = 8, hence middle one is 8 , the numbers is 7,8,9 same as others

OpenStudy (anonymous):

i did all but there is no answer

zepdrix (zepdrix):

Hmm, there is no correct answer D: that's so weird

OpenStudy (anonymous):

serious?

zepdrix (zepdrix):

18 = 5 + 6 + 7 24 = 7 + 8 + 9 33 = 10 + 11 + 12 36 = 11 + 12 + 13 Are you sure you posted the `options` correctly? :o